# The solving degrees for computing Gröbner bases of affine semi-regular polynomial sequences

### v2

@@ 4.1 Correspondence between $G_{\rm hom}$ and $G_{\rm top}$
Here we revisit the results in [30, Section 5.1].
Corollary 4.1.1 ([30, Corollary 2]).
@@
Since $\bm{F}^{\rm top}$ is cryptographic semi-regular and since $\bm{F}^{h}$ is $D$-regular by Corollary 3.1.5, we obtain $H_{1}(K_{\bullet}(F^{\rm top}))_{{\color[rgb]{0,0,0}<D}}=H_{1}(K_{\bullet}(F^{h}))_{{\color[rgb]{0,0,0}<D}}=0$. Moreover, as $H_{1}(K_{\bullet}(F^{h}))={\rm syz}(F^{h})/{\rm tsyz}(F^{h})$ and $H_{1}(K_{\bullet}(F^{\rm top}))={\rm syz}(F^{\rm top})/{\rm tsyz}(F^{h})$ (see (A.1.2)), we have the following corollary, where ${\rm tsyz}$ denotes the module of trivial syzygies (see Definition A.1.1).
Corollary 4.1.2 ([14, Theorem 1]).
@@
[…]
@@
We also note that the argument and the proof of Lemma 4.1.4 can be considered as a corrected version of [38, Theorem 4].
@@ 4.2 Correspondence between the computations of $G_{\rm hom}$ and $G$
In this subsection, we show that, at early stages, there is a strong correspondence between the computation of $G_{\rm hom}$ and that of $G$, from which we shall extend the upper bound on solving degree given in [40, Theorem 2.1] to our case.
@@
In [40], polynomial ideals over $R={\mathbb{F}}_{q}[x_{1},\ldots,x_{n}]$ are considered. Under the condition where the generating set $F$ contains the field equations $x_{i}^{q}-x_{i}$ for $1\leq i\leq n$, recall from Theorem 2.2.5 ([41, Theorem 6.5 & Corollary 3.67]) that the solving degree $\mathrm{sd}_{\prec}^{\mathcal{A}}(F)$ in the strict sense (see the definition (I) of Subsection 2.2 for the definition) with respect to a Buchberger-like algorithm $\mathcal{A}$ for $\langle F\rangle$ is upper-bounded by $22D-$ 2, where $D=d_{\rm deg}(\langle F^{\rm top}\rangle)$. In the proofs of [41, Theorem 6.5 & Corollary 3.67], the property $\langle F^{\rm top}\rangle_{D}=R_{D}$ was essentially used for obtaining the upper-bound. As the property also holds in our case, we may apply their arguments. Also in [4, Section 3.2], the case where $F^{h}$ is cryptographic semi-regular is considered. The results on the solving degree and the maximal degree of the Gröbner basis are heavily related to our results in this subsection.
@@
Here we show an upper-bound on the highest step degree appeared in the computation of $G$ with respect to the DRL ordering by a Buchberger-like algorithm $\mathcal{A}$ based on S-polynomials with the normal strategy and the sugar strategy. We note that, in [30, Lemma 4.2.4], we restart the computation of the Gröbner basis of $F$ from $H=\{g|_{y=1}:g\in(G_{\rm hom})_{\leq D}\}$. However, here we do not need $(G_{\rm hom})_{\leq D}$. We refer to [8, Remark 15] for another proof of ${\rm max.GB.}\deg_{\prec}(F)\leq D$.
Lemma 4.2.4 (cf. [30, Lemma 4]).
@@
We refer to [8, Remark 15] for another proof of ${\rm max.GB.}\deg_{\prec}({F})\leq D$. We also note that, if $D=d_{\rm reg}({F}^{\rm top})<\infty$, Lemma 4.2.3 and Lemma 4.2.4 hold without the assumption that ${F^{\rm top}}$ is cryptographic semi-regular.
Remark 4.2.6 (cf. [30, Section 5.2]).
@@
We review a simple example shown in [30, Example 1] and examine the correspondences discussed in this and the previous subsections.
@@ 4.3 Experimental observation and a variant of Fröberg’s conjecture
In this subsection, we observe actual behavior of Gröbner basis computation from a part of our experimental results. For experiments, we used Magma V2.25-3 [5]. In particular, the built-in functions Dimension, HilbertSeries, and GroebnerBasis were applied. In our experiments, given $n$, $m$, $(d_{1},\ldots,d_{m})$, and an odd prime $q$ with $n<m$ and $d_{i}\geq 2$ (for the case where $n=m$, see Remark 4.3.3 below), we generate inhomogeneous polynomials $f_{1},\ldots,f_{m}$ in $R=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ of degrees $d_{1},\ldots,d_{m}$ whose constant terms are all zero (as in Example 4.2.7), where coefficients are chosen uniformly at random. Note that, for any sequence $\bm{F}=(f_{1},\ldots,f_{m})$ generated in this way, the homogenization $F^{h}$ with $F=\{f_{1},\ldots,f_{m}\}$ always has a non-trivial root $(0,\ldots,0,1)$: Any sequence of $m$ polynomials (in $R$) that has at least one root is transformed by a linear coordinate change into a polynomial sequence $(f_{1},\ldots,f_{m})\in R^{m}$ such that each $f_{i}$ has no constant term. Moreover, $\bm{F}^{\rm top}=(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})$ is expected to be cryptographic semi-regular (in fact, semi-regular) with high probability. Therefore, our setting causes no loss of generality.
@@
${\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n+1}}\right]\right)}+1$ (4.3.1)
on the maximal Gröbner basis degree $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$, which is sharper than our upper-bound provided in Theorem 2 (1). This sharper bound comes from a property that $\bm{F}^{h}$ is generalized cryptographic semi-regular (i.e., $D^{\prime}$-regular), where we set $D^{\prime}:=\widetilde{d}_{\rm reg}(F^{h})$: We confirmed in our experiments that $\bm{F}^{h}$ satisfies this property in most cases. Here we use the following lemma (the proof is straightfoward):
@@
Let $\bm{F}=(f_{1},\ldots,f_{m})\in(R\smallsetminus K)^{m}$ be a sequence of not necessarily homogeneous polynomial. Assume that $n<m$. Then we have the following:
@@
    If ${F}^{h}$ has at least one solution and if $\bm{F}^{h}$ is $D^{\prime}$-regular (and thus $D^{\prime}<\infty$), then $D^{\prime}$ is upper-bounded by (4.3.1).
@@
    If $\bm{F}^{\rm top}$ is cryptographic semi-regular, then $D:=d_{\rm reg}(F^{\rm top})$ is also upper-bounded by (4.3.1) (this also holds even if $n=m$).
@@
$\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq\max\{D,D^{\prime}\}\leq{\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n+1}}\right]\right)}+1$
by Proposition 2.3.3. We compare exact values for several bounds, for $n=9$ and $10$ with some conditions on $m$ and $(d_{1},\ldots,d_{m})$ in Tables 14, where we set $D:=d_{\rm reg}(F^{\rm top})$ for the case where $\bm{F}^{\rm top}$ is cryptographic semi-regular.
Table 1: Exact values for several upper-bounds on $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ in the case where $n=9$, $n+1\leq m\leq 2n$, and $d_{1}=\cdots=\cdots=d_{m}=2$.
$m$ $10$ $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$
@@
Table 2: Exact values for several upper-bounds on $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ in the case where $n=10$, $n+1\leq m\leq 2n$, and $d_{1}=\cdots=d_{m}=2$.
$m$ $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$ $19$ $20$
@@
Table 3: Exact values for several upper-bounds on $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ in the case where $n=9$, $n+1\leq m\leq 2n$, $d_{1}=\cdots=d_{n}=3$, and $d_{n+1}=\cdots=d_{m}=2$.
$m$ $10$ $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$
@@
Table 4: Exact values for several upper-bounds on $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ in the case where $n=10$, $n+1\leq m\leq 2n$, $d_{1}=\cdots=d_{n}=3$, and $d_{n+1}=\cdots=d_{m}=2$.
$m$ $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$ $19$ $20$
@@
Note that, if $m=n+1$, then it follows from $\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n+1}}=\prod_{i=1}^{n+1}(1+z+\cdots+z^{d_{i}-1})$ that the bound (4.3.1) is equal to $\sum_{j=1}^{n+1}(d_{j}-1)+1$, which is equal to Lazard’s bound. As a more particular case, if $d_{i}=2$ for all $i$, then it is equal to $n+2$. On the other hand, recall from [4, Theorem 4.1] that $D=\lfloor(n+1)/2\rfloor+1$, and thus $12D-$ 1 is equal to $n+2$ if $n$ is odd and to $n+1$ if $n$ is even.
@@
If $m=n$, then it follows from $\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n}}=\prod_{i=1}^{n}(1+z+\cdots+z^{d_{i}-1})$ that $D=\sum_{j=1}^{n}(d_{j}-1)+1$, and we have
$\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n+1}}=\frac{\prod_{i=1}^{n}(1-z^{d_{i}})}{(1-z)^{n}}\cdot(1+z+z^{2}+\cdots),$
whose degree-$d$ coefficients are equal to a constant for $d\geq D-1$. In this case, the $D^{\prime}$-regularity of $F^{h}$ implies that $D^{\prime}\leq D$, whence $D^{\prime}\in\{D-1,D\}$ and $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq D$ by Proposition 2.3.3. Note that the equality $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})=D$ holds if $\langle\mathrm{LM}(\bm{F}^{h})\rangle$ is a weakly reverse lexicographic ideal.
@@
Let $K$ be an infinite field, and let $R=K[x_{1},\ldots,x_{n}]$. Let $d_{1},\ldots,d_{m}$ are integers larger than $1$, and let $f_{1},\ldots,f_{m}$ be polynomials in $R$ of degrees $d_{1},\ldots,d_{m}$ such that each $f_{i}$ has no constant term. Then, for given $K$, $n$, $m$, and $(d_{1},\ldots,d_{m})$, the property that $\bm{F}^{h}=(f_{1}^{h},\ldots,f_{m}^{h})$ is generalized cryptographic semi-regular (i.e., $D^{\prime}$-regular) is generic, where $D^{\prime}=\widetilde{d}_{\rm reg}(F^{h})$ is the generalized degree of regularity of $\bm{F}^{h}$ defined in Definition 2.3.1.
Conjecture 4.3.4 can be viewed as Fröberg’s conjecture [21]: A generic sequence of homogeneous polynomials is $D$-regular, where $D$ is the degree of regularity of the sequence.
@@
    As we described as above, the solving degree $\mathrm{sd}^{\rm mac}_{\prec^{h}}(F^{h})$ (which is equal to the solving degree $\mathrm{sd}^{\rm mac}_{\prec}(F)$ if $\prec$ is a DRL order, see Subsection 2.2) is also upper-bounded by $D^{\prime}$, which can be quite smaller than $12D-$ 1.
@@
    where $N_{F^{h}}$ is the number of projective roots of $F^{h}$ counted with multiplicity. This implies that the Hilbert driven algorithm can be effectively applied to the Gröbner basis computation of $F^{h}$, from which a Gröbner basis of $F$ is easily obtained.
@@
      $\langle\mathrm{LM}(\bm{F}^{h})\rangle$ is a weakly reverse lexicographic ideal. (Cf. Moreno-Socías conjecture [35].)
@@
The assumption that $\bm{F}^{h}$ is generalized cryptographic semi-regular (i.e., $D^{\prime}$-regular) could be useful to estimate the security of multivariate cryptosystems and signature schemes (or algebraic attacks based on Gröbner basis computation), see e.g., [22, Subsection 2.3], where the authors of [22] assume that the bound (4.3.1) gives a degree bound of the XL algorithm [11]. In fact, the bound (4.3.1) has been sometimes used in the cryptographic community without assuming the $D^{\prime}$-regularity of $\bm{F}^{h}$, see [22, Subsection 2.3] (in particular [22, Remark 1]) for details. Supposing the $D^{\prime}$-regularity of the homogenization of a target polynomial system, one could give a mathematically rigid explanation about the complexity of computing the Gröbner basis of the system.

### v3

@@ 4.1 Correspondence between $G_{\rm hom}$ and $G_{\rm top}$
Here we revisit the results in [31, Section 5.1].
Corollary 4.1.1 ([31, Corollary 2]).
@@
Since $\bm{F}^{\rm top}$ is cryptographic semi-regular and since $\bm{F}^{h}$ is $D$-regular by Corollary 3.1.4, we obtain $H_{1}(K_{\bullet}(F^{\rm top}))_{<D}=H_{1}(K_{\bullet}(F^{h}))_{<D}=0$. Moreover, as $H_{1}(K_{\bullet}(F^{h}))={\rm syz}(F^{h})/{\rm tsyz}(F^{h})$ and $H_{1}(K_{\bullet}(F^{\rm top}))={\rm syz}(F^{\rm top})/{\rm tsyz}(F^{h})$ (see (A.1.2)), we have the following corollary, where ${\rm tsyz}$ denotes the module of trivial syzygies (see Definition A.1.1).
Corollary 4.1.2 ([15, Theorem 1]).
@@
Corollary 4.1.2 implies that, in the Gröbner basis computation $G_{\rm hom}$ with respect to a graded ordering $\prec^{h}$, if an S-polynomial $S(g_{1},g_{2})=t_{1}g_{1}-t_{2}g_{2}$ of degree less than $D$ is reduced to 00, it comes from some trivial syzyzy, that is, $\sum_{i=1}^{m}(t_{1}a^{(1)}_{i}-t_{2}a^{(2)}_{i}-b_{i})\mathbf{e}_{i}$ belongs to ${\rm tsyz}(F^{h})_{<D}$, where $g_{1}=\sum_{i=1}^{m}a^{(1)}_{i}f_{i}^{h}$, $g_{2}=\sum_{i=1}^{m}a^{(2)}_{i}f_{i}^{h}$, and $S(g_{1},g_{2})=\sum_{i=1}^{m}b_{i}f_{i}^{h}$ is obtained by $\Sigma$-reduction in the $F_{5}$ algorithm (or its variant such as the matrix-$F_{5}$ algorithm) with the Schreyer ordering. Thus, since the $F_{5}$ algorithm (or its variant) automatically discards an S-polynomial whose signature is the LM of some trivial syzygy, we can avoid unnecessary S-polynomials. See [18] for the $F_{5}$ algorithm and its variant, and also for the syzygy criterion.
@@
We also note that the argument and the proof of Lemma 4.1.4 can be considered as a corrected version of [39, Theorem 4].
@@ 4.2 Correspondence between the computations of $G_{\rm hom}$ and $G$
In this subsection, we show that, at early stages, there is a strong correspondence between the computation of $G_{\rm hom}$ and that of $G$, from which we shall extend the upper bound on solving degree given in [41, Theorem 2.1] to our case.
@@
In [41], polynomial ideals over $R={\mathbb{F}}_{q}[x_{1},\ldots,x_{n}]$ are considered. Under the condition where the generating set $F$ contains the field equations $x_{i}^{q}-x_{i}$ for $1\leq i\leq n$, recall from Theorem 2.2.5 ([42, Theorem 6.5 & Corollary 3.67]) that the solving degree $\mathrm{sd}_{\prec}^{\mathcal{A}}(F)$ in the strict sense (see the definition (I) of Subsection 2.2 for the definition) with respect to a Buchberger-like algorithm $\mathcal{A}$ for $\langle F\rangle$ is upper-bounded by $22D-$ 2, where $D=d_{\rm deg}(\langle F^{\rm top}\rangle)$. In the proofs of [42, Theorem 6.5 & Corollary 3.67], the property $\langle F^{\rm top}\rangle_{D}=R_{D}$ was essentially used for obtaining the upper-bound. As the property also holds in our case, we may apply their arguments. Also in [5, Section 3.2], the case where $F^{h}$ is cryptographic semi-regular is considered. The results on the solving degree and the maximal degree of the Gröbner basis are heavily related to our results in this subsection.
@@
Here we show an upper-bound on the highest step degree appeared in the computation of $G$ with respect to the DRL ordering by a Buchberger-like algorithm $\mathcal{A}$ based on S-polynomials with the normal strategy and the sugar strategy. We note that, in [31, Lemma 4.2.4], we restart the computation of the Gröbner basis of $F$ from $H=\{g|_{y=1}:g\in(G_{\rm hom})_{\leq D}\}$. However, here we do not need $(G_{\rm hom})_{\leq D}$. We refer to [9, Remark 15] for another proof of ${\rm max.GB.}\deg_{\prec}(F)\leq D$.
Lemma 4.2.4 (cf. [31, Lemma 4]).
@@
We refer to [9, Remark 15] for another proof of ${\rm max.GB.}\deg_{\prec}({F})\leq D$. We also note that, if $D=d_{\rm reg}({F}^{\rm top})<\infty$, Lemma 4.2.3 and Lemma 4.2.4 hold without the assumption that ${F^{\rm top}}$ is cryptographic semi-regular.
Remark 4.2.6 (cf. [31, Section 5.2]).
@@
We review a simple example shown in [31, Example 1] and examine the correspondences discussed in this and the previous subsections.
@@ 4.3 Experimental observation and a variant of Fröberg’s conjecture
In this subsection, we observe actual behavior of Gröbner basis computation from a part of our experimental results. For experiments, we used Magma V2.25-3 [6]. In particular, the built-in functions Dimension, HilbertSeries, and GroebnerBasis were applied. In our experiments, given $n$, $m$, $(d_{1},\ldots,d_{m})$, and an odd prime $q$ with $n<m$ and $d_{i}\geq 2$ (for the case where $n=m$, see Remark 4.3.3 below), we generate inhomogeneous polynomials $f_{1},\ldots,f_{m}$ in $R=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ of degrees $d_{1},\ldots,d_{m}$ whose constant terms are all zero (as in Example 4.2.7), where coefficients are chosen uniformly at random. Note that, for any sequence $\bm{F}=(f_{1},\ldots,f_{m})$ generated in this way, the homogenization $F^{h}$ with $F=\{f_{1},\ldots,f_{m}\}$ always has a projective zero $(0:\cdots:0:1)$: Any sequence of $m$ polynomials (in $R$) that has at least one affine zero is transformed by a linear coordinate change into a polynomial sequence $(f_{1},\ldots,f_{m})\in R^{m}$ such that each $f_{i}$ has no constant term. Moreover, $\bm{F}^{\rm top}=(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})$ is expected to be cryptographic semi-regular (in fact, semi-regular) with high probability. Therefore, our setting causes no loss of generality.
@@
$D_{\rm new}:={\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n+1}}\right]\right)}+1$ (4.3.1)
on the maximal Gröbner basis degree $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$, which is sharper than our upper-bound provided in Theorem 2 (1). This sharper bound comes from a property that $\bm{F}^{h}$ is generalized cryptographic semi-regular (i.e., $D^{\prime}$-regular), where we set $D^{\prime}:=\widetilde{d}_{\rm reg}(\langle F^{h}\rangle_{R^{\prime}})$: We confirmed in our experiments that $\bm{F}^{h}$ satisfies this property in most cases. Here we use the following lemma (the proof is straightforward):
@@
Let $\bm{F}=(f_{1},\ldots,f_{m})\in(R\smallsetminus K)^{m}$ be a sequence of not necessarily homogeneous polynomials. Assume that $n<m$. Then we have the following:
@@
    If ${F}^{h}$ has at least one projective zero and if $\bm{F}^{h}$ is $D^{\prime}$-regular (and thus $D^{\prime}<\infty$), then $D^{\prime}$ is upper-bounded by $D_{\rm new}$ given in (4.3.1).
@@
    If $\bm{F}^{\rm top}$ is cryptographic semi-regular, then $D:=d_{\rm reg}(\langle F^{\rm top}\rangle_{R})$ is also upper-bounded by $D_{\rm new}$ given in (4.3.1) (this also holds even if $n=m$).
@@
$\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq\max\{D,D^{\prime}\}\leq D_{\rm new}$ (4.3.2)
by Proposition 2.3.6 (this holds even for an arbitrary graded monomial ordering on $R$). Therefore, the complexity of computing $G_{\rm hom}$ (and $G$) with respect to the number of arithmetic operations on $K$ can be estimated as
$O\left(m\binom{n+D_{\rm new}}{D_{\rm new}}^{\omega}\right)$ (4.3.3)
by Corollary A.4.2 and Remark A.4.4 below.
We also compare exact values for several bounds on $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$, for $n=9$ and $10$ with some conditions on $m$ and $(d_{1},\ldots,d_{m})$ in Tables 1 and 2, where we set $D:=d_{\rm reg}(\langle F^{\rm top}\rangle_{R})$ for the case where $\bm{F}^{\rm top}$ is cryptographic semi-regular, namely
$D={\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n}}\right]\right)}+1,$
which tends to be close to $D_{\rm new}$ for $m\gg n$.
Table 1: Exact values for several upper-bounds on $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ (which is equal to the solving degree of $F^{h}$ in this case) in the case where $n\in\{9,10\}$, $n+1\leq m\leq 2n$, and $d_{1}=\cdots=\cdots=d_{m}=2$. The first table is for the case $n=9$, and the second one is for the case $n=10$.
The number $m$ ($>n=9$) of polynomials $10$ $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$
@@
The number $m$ ($>n=10$) of polynomials $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$ $19$ $20$
@@
Table 2: Exact values for several upper-bounds on $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ (which is equal to the solving degree of $F^{h}$ in this case) in the case where $n\in\{9,10\}$, $n+1\leq m\leq 2n$, $d_{1}=\cdots=d_{n}=3$, and $d_{n+1}=\cdots=d_{m}=2$. The first table is for the case $n=9$, and the second one is for the case $n=10$.
The number $m$ ($>n=9$) of polynomials $10$ $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$
@@
The number $m$ ($>n=10$) of polynomials $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$ $19$ $20$
@@
Note that, if $m=n+1$, then it follows from $\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n+1}}=\prod_{i=1}^{n+1}(1+z+\cdots+z^{d_{i}-1})$ that the bound (4.3.1) is equal to $\sum_{j=1}^{n+1}(d_{j}-1)+1$, which is equal to Lazard’s bound. As a more particular case, if $d_{i}=2$ for all $i$, then it is equal to $n+2$. On the other hand, recall from [5, Theorem 4.1] that $D=\lfloor(n+1)/2\rfloor+1$, and thus $12D-$ 1 is equal to $n+2$ if $n$ is odd and to $n+1$ if $n$ is even. More precisely, assuming the $D^{\prime}$-regularity of $\bm{F}^{h}$, we have
${\rm HS}_{R^{\prime}/\langle F^{h}\rangle}(z)\equiv(1+z)^{n+1}\pmod{z^{D^{\prime}}},$
where $(1+z)^{n+1}=\sum_{i=0}^{n+1}\binom{n+1}{i}z^{i}$. In the expansion of $(1+z)^{n+1}$, the coefficient of $z^{(n+1)/2}$ is (resp. the coefficients of $z^{n/2}$ and $z^{(n+2)/2}$ are) maximal among the non-zero coefficients for an odd (resp. even) $n$, see Figure 1 for some specific $n$. In particular, if $n$ is even, then we have $D=n/2+1$ and $\dim_{K}(R^{\prime}/\langle F^{h}\rangle)_{D-1}=\dim_{K}(R^{\prime}/\langle F^{h}\rangle)_{D}$, which means that multiplication by-$y$ map $(R^{\prime}/\langle F^{h}\rangle)_{D-1}\longrightarrow(R^{\prime}/\langle F^{h}\rangle)_{D}$ is bijective. Thus, in the Gröbner basis computation of $F$, there is no degree fall at degree $D$ (in fact, up to $D$, see Subsections 4.1 and 4.2).
@@
If $m=n$, then we can prove $D^{\prime}=D-1$ and $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq D$ only assuming that $\bm{F}^{\rm top}$ is cryptographic semi-regular, and consequently $\bm{F}^{h}$ is generalized cryptographic semi-regular. Indeed, by [22, b) on page 121], we have the following coefficient-wise inequality:
${\rm HS}_{R^{\prime}/\langle F^{h}\rangle}(z)\geq\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n+1}}=\frac{\prod_{i=1}^{n}(1-z^{d_{i}})}{(1-z)^{n}}\cdot(1+z+z^{2}+\cdots).$ (4.3.4)
The degree-$d$ coefficients of the right hand side are equal to a constant for $d\geq D-1$, where $D=\sum_{j=1}^{n}(d_{j}-1)+1$ by $\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n}}=\prod_{i=1}^{n}(1+z+\cdots+z^{d_{i}-1})$. The $D$-regularity of $\bm{F}^{\rm top}$ implies that $\bm{F}^{h}$ is also $D$-regular, whence the degree $d$-coefficients of the both hand sides of (4.3.4) are equal to one another for each $d$ with $d\leq D-1$, so that we can easily check $D^{\prime}\geq D-1$ (see also the first assertion of Proposition 2.3.6). Moreover, recall from Remark 3.1.2 that ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(D-1)\geq{\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)$ for any $d\geq D-1$. Therefore, the inequality of (4.3.4) is in fact the equality, and hence $D^{\prime}=D-1$. Therefore, we obtain $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq D$ by Proposition 2.3.6, and also $\bm{F}^{h}$ is $D^{\prime}$-regular.
Note that the equality $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})=D$ holds if $\prec$ is a DRL ordering and if $\langle\mathrm{LM}(\langle{F}^{h}\rangle)\rangle$ is a weakly reverse lexicographic ideal, see Proposition 2.3.6.
Anyway, in the case where $n=m$, the complexity of computing $G_{\rm hom}$ (and $G$) with respect to the number of arithmetic operations on $K$ can be estimated as
$O\left(n\binom{n+D}{D}^{\omega}\right)$
with $D=\sum_{j=1}^{n}(d_{j}-1)+1$, by Corollary A.4.2 and Remark A.4.4 below.
Figure 1: The values of coefficients in $(1+z)^{n+1}$ for $(n,m)=(9,10)$ (the left figure) and $(n,m)=(10,11)$ (the right figure). The horizontal axis indicates the degree $i$ of $z^{i}$, and the vertical axis indicates the value of the coefficient of $z^{i}$. Note that $D=(n+3)/2$ for an odd $n$ and $D=(n+2)/2$ for even $n$, and thus $D-1=5$ for $n\in\{9,10\}$. See Remark 4.3.2 for a description.
@@
Let $K$ be an infinite field, and let $R=K[x_{1},\ldots,x_{n}]$. Let $d_{1},\ldots,d_{m}$ are integers larger than $1$, and let $f_{1},\ldots,f_{m}$ be polynomials in $R$ of degrees $d_{1},\ldots,d_{m}$ such that each $f_{i}$ has no constant term. Then, for given $K$, $n$, $m$, and $(d_{1},\ldots,d_{m})$, the property that $\bm{F}^{h}=(f_{1}^{h},\ldots,f_{m}^{h})$ is generalized cryptographic semi-regular (i.e., $D^{\prime}$-regular) is generic, where $D^{\prime}=\widetilde{d}_{\rm reg}(\langle F^{h}\rangle_{R^{\prime}})$ is the generalized degree of regularity of $\langle{F}^{h}\rangle_{R^{\prime}}$ defined in Definition 2.3.1.
Conjecture 4.3.4 can be viewed as a variant of Fröberg’s conjecture [22]: A generic sequence of homogeneous polynomials is $D$-regular, where $D$ is the degree of regularity of the ideal generated by the sequence.
Remark 4.3.5.
In Conjecture 4.3.4, the reason why we take all the constant terms of $f_{i}$’s to be zero (equivalently the ideal $\langle F\rangle_{R}$ vanishes at $(0,\ldots,0)$) is the following: For any inhomogeneous polynomial system having at least one affine zero $(a_{1},\ldots,a_{n})$, we can convert a system of inhomogeneous polynomials of the form $f_{1},\ldots,f_{m}$ as in the conjecture with an affine linear transformation sending $(a_{1},\ldots,a_{n})$ to $(0,\ldots,0)$.
@@
    As we described as above, the solving degree $\mathrm{sd}^{\rm mac}_{\prec^{h}}(F^{h})$ (which is equal to the solving degree $\mathrm{sd}^{\rm mac}_{\prec}(F)$ if $\prec$ is a DRL order, see Subsection 2.2) is also upper-bounded by $D_{\rm new}$ given in (4.3.1), which can be quite smaller than $12D-$ 1.
@@
    where $N_{F^{h}}$ is the number of projective zeros of $\langle F^{h}\rangle$ counted with multiplicity. This implies that the Hilbert driven algorithm can be effectively applied to the Gröbner basis computation of $F^{h}$, from which a Gröbner basis of $F$ is easily obtained.
@@
      $\langle\mathrm{LM}(\langle{F}^{h}\rangle)\rangle$ is a weakly reverse lexicographic ideal. (Cf. Moreno-Socías conjecture [36].)
@@
The assumption that $\bm{F}^{h}$ is generalized cryptographic semi-regular (i.e., $D^{\prime}$-regular) could be useful to estimate the security of multivariate cryptosystems (or algebraic attacks based on Gröbner basis computation), see e.g., [23, Subsection 2.3], where the authors of [23] assume that the bound (4.3.1) gives a degree bound of the XL algorithm [12]. In fact, the bound (4.3.1) has been sometimes used in the cryptographic community without assuming the $D^{\prime}$-regularity of $\bm{F}^{h}$, see [23, Subsection 2.3] (in particular [23, Remark 1]) for details. Supposing the $D^{\prime}$-regularity of the homogenization of a target polynomial system, one could give a mathematically rigid explanation about the complexity of computing the Gröbner basis of the system, and one can apply our estimation (4.3.3) for the complexity.